Find the mass of a thin funnel in the shape of a cone z = V x2 + y2,1 Szs 3 if its density function is p(x, y, z) = 8 - z.
Answer #1

z = 5 x 172 1723 density P(x, y, z) = 8-2 find mass mass : JJ P(X, Y, Z ) d V - 1 / (8-2) dv X=rcoso x2 + y2 72. = rsino 1 33